OpenStudy (brightemmerald):

Complete the square to solve the equation below. x^2 + 2x - 14 = 10 A. x = 5; x = -4 B. x = -6; x = 4 C. x = -1; x = 4 D. x = -1; x = 3

OpenStudy (dumbcow):

To complete the square : \[x^2 +bx = c\] \[x^2 +bx + (\frac{b}{2})^2 = c + (\frac{b}{2})^2\] \[(x+ \frac{b}{2})^2 = c+ (\frac{b}{2})^2\] \[x + \frac{b}{2} =\pm \sqrt{c+ (\frac{b}{2})^2}\] \[x =- \frac{b}{2} \pm \sqrt{c+ (\frac{b}{2})^2}\]

OpenStudy (dumbcow):

In this example, b = 2 c = 10+14 = 24